A particle has a net charge of +0.31 nC.
How many electrons must be added to this particle in order to make it electrically neutral? (to 2 s.f)
1
Expert's answer
2020-05-05T18:41:58-0400

Solution. It is necessary to add an electric charge q= -0.31 nC to make a particle electrically neutral.

Since the charge of an electron qe= -1.6 x 10^-19C required number of electrons is equal to

"N=\frac{q}{q_e}=\frac{-0.31 \times 10^{-9}C}{-1.6 \times 10^{-19}C}\approx 1.94 \times 10^9"

Answer.

"N=1.94 \times 10^9"